silvernight269:

Simplify log_64 32 = x

silvernight269:

@darkknight

darkknight:

Log base 64 of 32?

silvernight269:

i think?

darkknight:

So log equations are written in this form: log x = y b Then b^y = x

darkknight:

So in this case Log 32 = x 64 64^x = 32 Can you solve for x?

silvernight269:

1/6?

darkknight:

Well no, 64^(1/6) would be 2 But think about what it would be. If you do take the 6th root, then do to the power of 5. It will be the same as 5/6, and that would get you to 32

darkknight:

So solving for x would get us 5/6

darkknight:

I don't think that was a good explanation, so you have 64^x = 32. Are you asking how to solve that?

silvernight269:

yes

silvernight269:

I dont really get it until i show through work lol

darkknight:

Okay, so how to do this to take the log base of both sides. So here we have 64^x = 32 so we take \[\log_{64} \] of both sides to cancel out the 64. So x = \[\log_{64}(32) \]

silvernight269:

ok so i got that

darkknight:

Alright. So if you ever approach a problem like that and have a variable as the exponent, what you can do is to take the log base of the base. So for ex: If you have b^y = x And y is your variable, then you can preform this: \[\log_{b}(b)^y = \log_{b}x \] which equals \[y= \log_{b}x \] Which, rewriting using change of base formula, is \[y = \log(x)/\log(b)\]

darkknight:

Hope that helped :)